18. Mohan and Sohan solve an equation. In solving Mohan commits a mistake in constant term and finds the roots 8 and 2. Sohan commits a mistake in the coefficient of x. The correct roots are
(a) 9,1
(b) -9,1
(c) 9, -1
(d) -9, -1

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: a Explaination:Reason: Correct sum = 8 + 2 = 10 from Mohan Correct product = -9 x -1 = 9 from Sohan ∴ x² – (10)x + 9 = 0 ⇒ x² – 10x + 9 = 0 ⇒ x² – 9x – x + 9 ⇒ x(x – 9) – 1(x – 9) = 0 ⇒ (x-9) (x-l) = 0 . ⇒ Correct roots are 9 and 1